# Active Alignments of Lens Systems with Reinforcement Learning*

### v1

@@ IV-C Generating the Reference Pattern
To decide whether a state $s$ is in the optimality condition, the observation $O_{w}(s)$ has to be compared to a reference pattern $O^{*}$. This reference pattern only depends on the number of lenses $k$ in the lens system, not on the noise level of $W$. To generate $W$ in our synthetic setting, we use a perfectly aligned lens system where each of the $k$ many lenses are perfectly aligned as well, i.e., $W^{*}:=0$. We then sample $1.000$ observations from $O_{W^{*}}(s^{*})$ and compute the pixel-wise mean image as an approximation for $O^{*}=\mathbb{E}[O_{W^{*}}(s^{*})]$.
V Experiments
V-A Benchmark Environments
@@
When rendering images, one typically faces a trade-off between computational efficiency and image quality. To reduce overall RL training time, we prioritize fast image generation at the cost of lower pixel resolution and increased sensor variance. Because of that, we only use $50\times 50$ pixels and $512$ samples per pixel for rendering. This also motivates the decision to use a simple rectangle as reference pattern, as more detailed patterns are not recognizable at this resolution.
@@
For the evaluation, we focus on six distinct benchmark setups. Each setup considers a set of $k\in\{2,3\}$ single lenses with radii of curvature of $-1.0$ and $3.8$ and thickness of $0.2$. Each single lens can have none, low or high individual variances $W_{L}$. The first three components of $W_{L}$ represent translation offsets along the $x$, $y$ and $z$ axes and are sampled from a normal distribution $(W_{L})_{i}\sim N_{0,6.25}$ for low and $(W_{L})_{i}\sim N_{0,12.5}$ for high variances with $i\in\{1,2,3\}$. The remaining two represent rotation offsets along the $x$ and $y$ axes, sampled as $(W_{L})_{j}\sim N_{0,0.00625}$ for low and $(W_{L})_{j}\sim N_{0,0.0125}$ for high variances with $j\in\{4,5\}$. These variances account for potential manufacturing imperfections within a certain tolerance bound. To create a challenging alignment situation, they are intentionally set very high (see Figure 2 for an extreme case). Furthermore, rotation around $z$ is considered redundant, as in a perfectly aligned scenario, rotation around the $z$ axis has no effect. Note that $O^{*}$ does not change for setups including object variances.
V-B Algorithms
Expert-designed alignment algorithms typically rely on hand-crafted features extracted from $O_{W}(s)$ that need to be optimized simultaneously. These features as well as their optimality region highly depend on the exact use case. To have a fair and generic comparison of our RL approach with conventional learning methods, we decided to benchmark against state-of-the art black-box optimizers to solve Equation (1) actively. For our study, we focused on three different models to solve the task of optical alignments:
V-B1 Proximal Policy Optimization
Proximal policy optimization (PPO) is a policy-based RL algorithm that trains a stochastic policy in an on-policy way [26]. To process the images, a CNN is used as policy network. While PPO is a versatile learning algorithm, many learning tasks are too complex to solve from scratch via RL. Instead of training an agent on a hard task from the start, curriculum learning structures the training by introducing simpler subtasks first, enabling the agent to build foundational skills before tackling more challenging scenarios. This typically leads to faster convergence, better generalization, and improved sample efficiency, especially in environments with sparse rewards or long-horizon decision-making problems [27]. The same holds true when training an RL agent for a too low optimality threshold $\theta$: The agent fails to reach the optimality region, even when allowing long trajectory lengths before truncation. Thus, we apply a curriculum strategy where $\theta$ is decreased gradually in an exponential fashion, that is, starting from an initial threshold $\theta_{0}$, the threshold of the $i$-th task is $\theta_{i}=C\cdot\theta_{i-1}$ with $C<1$ (in our experiments, we set $C=0.9$). We denote the task as solved when the agent manages to reach the optimality region defined by $\theta_{i}$ within $20$ steps in five subsequent online evaluations. The maximal length of an episode $T$ is set to $100$. All models have been trained for $1.5\cdot 10^{6}$ global steps with an initial learning rate of $1e^{-4}$ which is decreased by a factor of $0.8$ if the reward stagnates and which is reset to the initial value if the task of the curriculum is solved. Figure 4 shows the evolution of the reward along the curriculum for $k=3$ lenses. All trainings are executed on NVIDIA H100 GPUs requiring roughly $0.32$ seconds per global step resulting in a total train time of approximately five days.
V-B2 Bayesian Optimization
@@
Bayesian Optimization (BO) is a strategy for minimizing functions $f$ that are expensive to evaluate by building a probabilistic model of the objective function [28, 29]. It selects new evaluation points by balancing exploration and exploitation using an acquisition function, which predicts where the function is likely to improve. This approach is particularly useful when function evaluations are costly, as it finds optimal solutions with relatively few evaluations. Setting $f_{W}(s):=\|O_{W}(s)-O^{*}\|$, an alignment problem can be interpreted as a black-box optimization problem. We evaluated BO algorithms using a different probabilistic model than implemented in scikit-optimize [30]: Gaussian Processes (BO-GP) [31] and Random Forests (BO-RF) [32]. As vanilla BO algorithms explore $f_{W}$ for each $W$ and without a priori information, the state space has to be explored first randomly costing unnecessary steps. Thus, we also tested the method proposed in [33] (TransferGP) that allows pre-training a Gaussian process on samples from problem instances $f_{W_{1}},\ldots,f_{W_{m}}$.
V-B3 Random
@@
V-C Results
For evaluation, each algorithm was executed on $100$ different environments for each of the six benchmark situations as described in Section V-A. We compared approaches that operate without any a priori information and require no domain knowledge. Except for the RL algorithms, the algorithms used do not involve a training phase. To address this imbalance, we reduced the search space for baseline algorithms to approximately eight percent of the search space used for RL. Figure 5 shows the RL-based method trained with $R_{\mathrm{pot}}$ surpases in all scenarios all other algorithms in terms of convergence speed. Moreover, independent of the number of lenses and the noise level, RL-PPO reaches the optimality region in under ten steps. As expected, the benchmarks with larger noise levels converge at a higher error, because due to the lens variances, even the best alignment cannot reach the minimal render variance error. In a direct comparison, the potential-based reward $R_{\mathrm{pot}}$ converges slighly faster than $R_{\mathrm{step}}$ to the optimal region (see Figure 6).
Figure 4: Achieved reward $R_{\mathrm{pot}}$ during an online evaluation for $k=3$ lenses and different noise levels while optimality threshold $\theta$ is adjusted according to the curriculum.
Only considering the computation time, RL-PPO requires a constant 25 ms and BO-RF 95 ms per alignment step. In contrast, BO-GP becomes increasingly time consuming as the number of steps increases. For 20 steps, each step takes an average of 100 ms, while for 50 steps the time per step increases to 178 ms. For the pre-trained Gaussian processes of TransferGP, the processing time for each step increases with the number of instances it has been pre-trained on. Already when trained on $m=10$ instances with $100$ samples each, their processing time per step takes several minutes while their performance equals almost the performance of vanilla BO-GP. Due to their impractical computation time, we have not included pretrained GP models in our benchmark.
Figure 5: Comparison of alignment algorithms.
Figure 6: Evaluation of RMSE for the best PPO models. The upper plot compares two reward functions for a fixed number of single lenses (three). The lower plot compares two different lens systems using the best reward function.
VI Conclusion
@@
ACKNOWLEDGMENT
We thank Marco F. Huber and our colleagues from the Institute of Machine Vision for their careful proofreading and constructive feedback on the manuscript.

### v2

@@ IV-C Generating the Reference Pattern
To decide whether a state $s$ is in the optimality condition, the observation $O(s,W)$ has to be compared to a reference pattern $O^{*}$. This reference pattern only depends on the optical layout of the lens system, not on the noise level of $W$. To generate $W$ in our synthetic setting, we use a perfectly aligned lens system where each of the single lenses are perfectly aligned as well, i.e., $W^{*}:=0$. We then sample $1.000$ observations from $O_{W^{*}}(s^{*})$ and compute the pixel-wise mean image as an approximation for $O^{*}=\mathbb{E}[O_{W^{*}}(s^{*})]$.
V Measuring the sim-to-real gap
To emulate realistic manufacturing scenarios, we validate our simulation relign against an established optical design software and images from a real-world alignment station. For all the results presented in Section VI, we used the following hand-designed lens system was used.
V-A Comparison to Optical Design Software
To assess the applicability of the proposed workflow to a more realistic optical system, we designed a lens system consisting of four lenses in Code V [26] for monochromatic operation at 589 nm. The system is optimized for imaging objects at infinity onto the sensor plane and exhibits an effective focal length (EFL) of 2.895 mm, an f-number of 2.1, a field of view (FOV) of 84.1°, and a half exit-pupil angle of 13.45°. The lens system has a mechanical length of 20.59 mm and a maximum diameter of 9.8 mm. The design introduces a minimal distortion of –27.2% and throughout the entire field of view, the relative illumination does not fall below 60%. Figure 3 shows an optical diagram of the lens system including light rays of different wave lengths.
Figure 3: Optical diagram of lens system including light rays in Code V.
For integration into relign, each lens has to be exported individually from Code V and converted into polygonal meshes using any CAD tool (e.g., SolidWorks). Exporting the elements separately, rather than the lens system as a whole, ensures that the components remain independently addressable within the simulation environment. To enforce the defined field of view, we introduced an additional mechanical mount and an aperture stop directly behind the first lens.
We modeled illumination by a rectangular area emitter providing diffuse emission with arbitrary bitmap textures. The emitter dimensions and distance were chosen such that the system operates within the optical field of view and the incident rays approximate collimated illumination, consistent with the design specifications.
Simulation accuracy was validated using a chessboard test pattern. Subpixel-accurate corner positions were extracted from both the Code V and Mitsuba sensor outputs. The mean positional deviation amounted to approximately 1.2 pixels in both x- and y-directions, indicating close agreement between design and simulation as can be seen in Figure 4.
(a) relign (ours)
(b) Code V
(c) Corner detection differences
Figure 4: Comparison of point matching on a chessboard grid: relign (ours) vs. Code V.
V-B Comparison to Real Active Alignment Situation
Established or publicly documented alignment methods for complex multi-lens systems are rare hindering a realistic and unbiased comparison of different active alignment methods. This is mostly due to strong confidentiality constraints surrounding industrial active alignment, which typically cover algorithms, optical designs, hardware, and in some cases even the alignment targets.
Consequently, we designed along given the optical parameters of a real-world camera system an optical design that approximates the performance. We then compare the output of relign to a real sensor output of $1920\times 1280$ pixel obtained through an alignment station. The visual comparison is shown in Figure 5 demonstrating that relign closely approximates a realistic scenario, with only a small deviation from the real-world camera.
(a) relign (ours)
(b) Real image from production
(c) Difference relign (red) and production (green)
Figure 5: Comparison of a realistic alignment target: relign (ours) vs. a real alignment station.
VI Experiments
We conduct a benchmark of different active alignment methods in this section to showcase how relign can be used. Throughout, we use the optical layout introduced in Section V.
@@ VI Experiments
When rendering images, one typically faces a trade-off between computational efficiency and image quality. To reduce overall RL training time, we prioritize fast image generation at the cost of lower pixel resolution and increased sensor variance. Because of that, we only use $200\times 200$ pixels and $64$ samples per pixel for rendering.
Accurate and comparable benchmarks for active alignments are hard to find, not only due to privacy constraints of companies maintaining the algorithms and the optical designs as part of their core know-how, but also due to the fact that optical products are very diverse and have different requirements. Already small differences in the required optical performance or the assumed variances of the optical system can lead to completely different alignment strategies and performances. Typical expert-designed alignment algorithms rely on hand-crafted features extracted from $O(s,W)$ that need to be optimized simultaneously, often via curve fitting [27] or expensive area scans. Already for use cases with two degrees of freedoms, dozens of steps may be required [27].
A comparison of RL algorithms with classic methods on real systems thus requires a careful design to make them truly unbiased and insightful and hence this is beyond the scope of this paper. Nevertheless, we are convinced that our framework eases the effort to design fair and direct comparison in the future. Instead, we decided to benchmark against state-of-the art optimizers to solve Equation (1) actively and ensure that all algorithms are confronted with the exact same problem instances.
VI-A Benchmark Environments
@@
For the evaluation, we focus on three distinct benchmark setups. Each setup considers the same lens system, but with either none, low or high individual variances $W_{L}$ for the last two lenses. The first three components of $W_{L}$ represent translation offsets along the $x$, $y$ and $z$ axes and are sampled from a normal distribution $(W_{L})_{i}\sim N_{0,1.25\cdot 10^{-4}}$ for low and $(W_{L})_{i}\sim N_{0,2.5\cdot 10^{-4}}$ for high variances with $i\in\{1,2,3\}$. The remaining two represent rotation offsets along the $x$ and $y$ axes, sampled as $(W_{L})_{j}\sim N_{0,0.375}$ for low and $(W_{L})_{j}\sim N_{0,0.75}$ for high variances with $j\in\{4,5\}$. We denote the underlying distributions as $\rho_{N}$ for $N=0$, $N=0.25$ and $N=0.5$. These variances account for potential manufacturing imperfections within a certain tolerance bound. To create a challenging alignment situation, they are intentionally set very high. Furthermore, rotation around $z$ is considered redundant, as in a perfectly aligned scenario, rotation around the $z$ axis has no effect. Note that $O^{*}$ does not change for setups including object variances.
VI-A1 RL Algorithms
Proximal policy optimization (PPO) is a policy-based RL algorithm that trains a stochastic policy in an on-policy way [28].
We used the PPO implementation in Stable Baselines3 [29] and compared it with other well-tested State-of-the-art RL algorithms. Our analysis revealed PPO’s superior performance, leading us to adopt it as the exclusive RL approach for this study.
To process image observations, we employed a CNN as policy network, that consists of a standard CNN architecture with three $3\times 3$ convolutional layers, each followed by ReLU activation and max-pooling. This is followed by a single fully connected layer holding $256$ extracted features for each image.
As is common in previous work, the learning rate and the entropy coefficient are crucial hyperpameters of PPO. We experimented with different learning rates in $\{1e^{-2},1e^{-3},1e^{-4},1e^{-5}\}$ and entropy coefficients in $\{1e^{-2},1e^{-3},1e^{-4}\}$.
We set the discount factor to $\gamma=0.9$ and used the reward function described in Section III with $C=0.1$. The best-performing models were obtained after $1.9\cdot 10^{6}$ global steps and learning rates of $10^{-3}$ and $10^{-4}$, dependent on $W_{L}$.
Figure 6 shows the evolution of the reward during online evaluations. Counterintuitively, one would expect to see a much slower convergence for higher noise during training. But due to the high variance in high-noise scenarios, the globally defined threshold $\theta$ must be set at a higher level to ensure that even lower-quality lens system still meet the acceptance criteria. Otherwise, such lens systems would be classified as not acceptable. However, increasing the tolerance in this way inevitably leads to a reduction in overall manufacturing quality.
All trainings are executed on NVIDIA L40S GPUs requiring roughly $0.05$ seconds per global step resulting in a total train time of approximately $28$ hours. The full details can be found in the accompanying code of this paper.
VI-A2 Bayesian Optimization
@@
Bayesian Optimization (BO) is a strategy for minimizing functions $f$ that are expensive to evaluate by building a probabilistic model of the objective function [30, 31]. It selects new evaluation points by balancing exploration and exploitation using an acquisition function, which predicts where the function is likely to improve. This approach is particularly useful when function evaluations are costly, as it finds optimal solutions with relatively few evaluations. Setting $f_{W}(s):=-R(s,W)$, an alignment problem can be interpreted as an optimization problem for the family of black box functions $f_{W}$. We evaluated BO algorithms using a different probabilistic model than implemented in scikit-optimize [32]: Gaussian Processes (BO-GP) [33] and Random Forests (BO-RF) [34]. As vanilla BO algorithms explore $f_{W}$ for each $W$ and without a priori information, the state space has to be explored first randomly costing unnecessary steps. Thus, we also tested the method proposed in [35] (TransferGP) that allows pre-training a Gaussian process on samples from problem instances $f_{W_{1}},\ldots,f_{W_{m}}$.
VI-A3 Random
@@
VI-B Results
For evaluation, each algorithm was executed on $100$ different environments for each of the benchmark situation described in Section VI-A. We compared approaches that operate without any a priori information and require no domain knowledge. Except for the RL algorithms, the algorithms used do not involve a training phase. To address this imbalance, we reduced the search space for baseline algorithms to approximately eight percent of the search space used for RL. The root mean squared error (RMSE), computed over all pixels, was used as the performance metric. Figure 7 shows that the RL-based method surpasses in all scenarios all other algorithms in terms of convergence speed. Moreover, independent of the noise level, RL-PPO reaches its optimum in under ten steps. As expected, the benchmarks with larger noise levels converge at a higher error, because due to the lens variances, even the best alignment cannot reach the minimal render variance error.
Figure 6: Evolution of rewards during training for different noise levels.
Only considering the computation time, RL-PPO requires a constant 25 ms and BO-RF 95 ms per alignment step. In contrast, BO-GP becomes increasingly time-consuming as the number of steps increases. For 20 steps, each step takes an average of 100 ms, while for 50 steps the time per step increases to 178 ms. For the pre-trained Gaussian processes of TransferGP, the processing time for each step increases with the number of instances it has been pre-trained on. Already when trained on $m=10$ instances with $100$ samples each, their processing time per step takes several minutes while their performance equals almost the performance of vanilla BO-GP. Due to their impractical computation time, we have not included pre-trained GP models in our benchmark.
Figure 7: Comparison of baseline alignment algorithms against RL-PPO for different noise levels.
Figure 8: Evaluation of RMSE for the best PPO models and different noise levels.
VII Conclusion